In R’s built-in datasets library is the data frame quakes. Make sure you can access this object. Use the command head(quakes).

Then, do the following:

Select only those records that correspond to a magnitude (mag) of greater than or equal to 5 and write them to a file called Lab4-Q2.txt

You can use the command

dput(x=YourDataVariable, file="fileName.txt")
Answer #1

Herethe library tidyerse is used.

library(tidyverse)

mag_quakes<-quakes %>% He
filter(mag >= 5.0)

load(mag_quakes)

" Here the dput is used and the output can be on R-studio console with the first 5 rows"

dput(mag_quakes[1:5,c(1:5)])
structure(list(lat = c(-26.00, -20.70, -13.64, -19.66, -16.46),
long = c(184.10, 169.92, 165.96, 180.28, 180.79),
depth = c(42, 139, 50, 431, 498),
mag = c(5.4, 6.1, 6.0, 5.4, 5.2),
stations = c(43, 94, 83, 57, 79)), .Names = c("lat","long","depth","mag","stations"), row.names = c(NA, 5L), class = "data.frame")

"Here the output is send to a file"
dput(mag_quakes,"Lab4-Q2.txt")